### v3

@@ 5.3 Anchoring and Calibration Procedures
This paper distinguishes between two steps in adapting a scoring model to operational use:
  • Anchoring defines the maximum acceptable error limit for a standard amount of text. This tolerance is expressed in penalty points based on a user’s real-world standard, such as, “no more than 5 penalty points per 1,000 words.” (one minor error is typically equivalent to one penalty point).
  • Calibration takes the absolute penalty point total and maps it onto a normalized, more easily interpretable 0-100 quality scale for use in dashboards or QA reports.
Feasibility (two‑point).
Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
@@
Intuitively, the longer sample must permit more errors but sublinearly; the shorter must permit fewer errors but not in direct proportion to length.
The model can be easily calibrated using two known tolerance points. Suppose we have:
  • The primary anchor point: This is the main tolerance defined during the Anchoring step, such as a maximum of 5 errors ($E_{0}=5$) for a reference sample size of 1,000 words ($x_{0}=1000$).
  • A second, empirical data point: This point is derived from observing the user’s real-world judgment (e.g., through a survey), such as their intuitive tolerance for a different sample size—for instance, accepting only 2 errors ($E_{1}=2$) in a 250-word text ($x_{1}=250$).
Two‑point calibration (numerical).
Let $r=E_{1}/E_{0}$ and define $f(b)=\ln(1+bx_{1})-r\,\ln(1+bx_{0})$. Under the feasibility condition above, $f$ has exactly one root $b^{\star}>0$ by the intermediate value theorem. Near $b=0$, $f(b)\approx b(x_{1}-rx_{0})$, so the sign at the origin is $\mathrm{sgn}(x_{1}-rx_{0})$. As $b\to\infty$, $f(b)\sim(1-r)\ln b+\ln x_{1}-r\ln x_{0}$; if $r>1$ the limit is -$\infty$, and if $r<1$ it is $\infty$, guaranteeing a sign change and thus a unique root. Solve $f(b)=0$ with a robust one‑dimensional method (e.g., bisection/Brent); then set
@@
See  Calibration from Two Points (Numerical) for a one-dimensional root-finding procedure in $b$ (e.g., bisection), followed by $a=E_{0}/\ln(1+b\,x_{0})$.
@@
Feasibility.
Ensure $r=E_{1}/E_{0}$ lies strictly between $1$ and $\rho=x_{1}/x_{0}$, i.e., $\min\{1,\rho\}<r<\max\{1,\rho\}$, before solving for $b$.
Solving $f(b)=0$ yields $b^{\star}\approx 2.880\times 10^{-3}$ and $a^{\star}\approx 3.688$, giving the calibrated curve
@@
Fallback for multiple points. If more than two tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ are available—or if the two chosen points violate the feasibility condition—estimate $(a,b)$ by constrained least squares:
@@
Given any $b_{0}>0$, a fast 1‑D search over $b$ with the closed‑form $a(b)=\dfrac{\sum_{i}E_{i}\,\ln(1+bx_{i})}{\sum_{i}\ln^{2}(1+bx_{i})}$ works well (then refine both jointly if desired).
This approach allows the model to be tailored to specific content types, organizational standards, or language pairs while still operating under a common mathematical framework. For instance, the error tolerance can be precisely defined for each case, ranging from extremely strict for software UI text, to moderately strict for marketing copy and documentation, and significantly more lenient for internal communications or user-generated content.
For small $b$, $\ln(1+b\,x)\approx b\,x$, so the model reduces locally to a linear tolerance rule; this explains why a calibrated linear slope performs acceptably only in the vicinity of its reference sample size.
If $x$ is expressed in pages with $W$ words per page, the same curve is obtained by replacing $b$ with $W\cdot b$; $a$ is unchanged.
Figure 4: Three interpolations. The spreadsheet curve $c+k\ln x$ is shown for comparison only; calibration uses $E(x)=a\ln(1+bx)$, which passes through the origin and is anchorable to tolerance points.
5.4 Near-Reference Fidelity Interval for Linear Approximation (±20%)
In practice, a linear rule anchored at a reference size $x_{\mathrm{ref}}$ is often used for convenience:
@@ 5.4 Near-Reference Fidelity Interval for Linear Approximation (±20%)
To quantify where this linear approximation is acceptable, we define the ±20% fidelity interval around $x_{\mathrm{ref}}$ as the set of $x>0$ such that
@@
Writing $\varepsilon=0.20$ and $\alpha=\dfrac{\ln(1+b\,x_{\mathrm{ref}})}{b\,x_{\mathrm{ref}}}$, the two boundary points admit a closed form via the Derivation (sketch). Set $E_{\rm lin}(x)=(x/x_{\rm ref})\,E_{\rm log}(x_{\rm ref})$ and require $E_{\rm lin}(x)=(1\pm\varepsilon)E_{\rm log}(x)$. Writing $\alpha=\ln(1+bx_{\rm ref})/(bx_{\rm ref})$ and $y=1+bx$, we obtain $\alpha\ln y=(1\pm\varepsilon)(y-1)/y$. Rearranging gives $y\,e^{-\frac{1\pm\varepsilon}{\alpha}\,y}=e^{-\frac{1\pm\varepsilon}{\alpha}}$, so
$y=-\frac{1\pm\varepsilon}{\alpha}\,W_{-1}\!\Bigl(-\frac{\alpha}{1\pm\varepsilon}\,e^{-\alpha/(1\pm\varepsilon)}\Bigr),$
which yields the closed form in the next equation. The branch $W_{-1}$ selects the solution near $x_{\rm ref}$.
Lambert $W$ function:
@@
where $x_{-}$ (with the “-” sign) gives the lower boundary (linear is $0.8\times$ the log curve) and $x_{+}$ (with the “++” sign) gives the upper boundary (linear is $1.2\times$ the log curve). [^8] The branch $W_{-1}$ is selected to obtain the solution near $x_{\mathrm{ref}}$.
Let’s look at the numerical intervals for the example calibration. Using the illustrative curve from §5.3 (a=3.688,b=0.00288a=3.688,\,b=0.00288), the ±20% fidelity intervals are:
  • $\boldsymbol{x_{\mathrm{ref}}=1000}$ words: $x\in[579,\;1460]$ words, width $\approx 881$ words (about $42\%$ to $46\%$ relative to $x_{\mathrm{ref}}$).
  • $\boldsymbol{x_{\mathrm{ref}}=2000}$ words: $x\in[1307,\;2747]$ words, width $\approx 1440$ words (about $34.7\%$ to $37.4\%$ relative to $x_{\mathrm{ref}}$).
These calculations illustrate how far one can move away from the reference sample size while still retaining acceptable measurement accuracy. The window is relatively narrow: if the reference is 1,000 words, then evaluating a 2,000-word sample already falls outside the ±20% fidelity band and will not be accurate; likewise, if the reference is 2,000 words, applying the same linear rule to a 1,000-word sample is equally unreliable.
In absolute terms, the ±20% window widens as the reference size increases (from about 0.88 k to about 1.44 k words), but in relative terms it becomes proportionally narrower. This quantifies the earlier point: a calibrated linear rule is only reliable near its anchoring point. Further away, it becomes too harsh for short samples and too lenient for long ones. When jobs regularly fall outside these fidelity intervals, the logarithmic model should be used directly, or the linear rule should be re-anchored to a more suitable reference size.
5.5 Visual comparison across three sample-size ranges
@@
Figure 5: Linear model (blue) vs. human perception (PFT, red). Yellow bands indicate the $\pm 20\%$ fidelity zone around the linear anchor; outside these bands the linear rule over‑ or under‑estimates true tolerance.
Figure 5 visualizes this critical weakness of the linear model: Once the slope is fixed using a single calibration point, the linear model severely diverges from the actual human judgment at both lower and higher evaluation sample sizes. In contrast, the logarithmic model aligns more closely with observed data regardless of sample size (evaluation word-count).
The calibrated logarithmic curve is more permissive for short samples and stricter for long ones relative to an anchored linear rule; it increases with length but flattens, avoiding the linear model’s over‑leniency at large sizes. This not only matches human content users’ expectations, but also reduces misclassification of borderline quality outputs.
@@
Figure 6: Applicability across evaluation sample sizes. SQC (micro, $\lesssim\!250$ words), calibrated linear MQM (meso; reliable only within $\sim\!\pm 20\%$ of anchor), and non-linear MQM (macro). Vertical mark at $\sim\!250$ words flags the micro‑range where deterministic curves are unreliable. Y‑axis: error tolerance (penalty points).
In summary, building on the Multi-Range Framework [GladkoffEtAl2024], this paper distinguishes three regimes of evaluation sample sizes based on their statistical and perceptual characteristics. In the micro-range (fewer than 250 words), evaluation outcomes are highly sensitive to single errors and statistical confidence is very low; thus, risk-based Statistical Quality Control (SQC) is most appropriate.
For the meso- and macro-ranges, the choice of an appropriate scoring model depends not on the absolute length of any single text, but on the expected range of fluctuation in evaluation sample sizes. This paper’s findings point to a crucial rule of thumb: if evaluation sample sizes are expected to vary by more than approximately 20% from the anchor point, the non-linear model is required for accuracy. A calibrated linear model is only a viable tool when evaluation sample sizes remain consistently within this tight  20% band around the anchor. Counter-intuitively, because the logarithmic curve flattens at large values, a linear model can be more applicable in the macro-range than the meso-range, provided its anchor point is also in that higher range.
Ultimately, however, for any operational scenario involving a wide variety of evaluated text lengths, the non-linear logarithmic model is the only option that ensures consistent and reliable accuracy. This is because a single linear extrapolation becomes unreliable as it moves further from its anchor point, whereas a logarithmic model remains accurate across both the meso- and macro-ranges. Figure 6 summarizes these applicability ranges, illustrating how tolerance curves change shape across evaluation sample size regimes.
@@
EWC varies within $\pm 20\%$ around the anchor Calibrated linear model (re-anchor if the range shifts)
@@
Very short samples ($\lesssim 250$ words) Statistical Quality Control (SQC); deterministic curves unreliable
Table 1: Operational guidance: which scoring model to use in practice.
@@ 6 Practical Implications
To the extent that error tolerance is not linear — as suggested by both empirical findings and practitioner reports — the adoption of a logarithmic scoring model can reshape translation quality evaluation practices, quality management processes, and software tooling across the industry. Replacing the linearity assumption allows quality assessment to become more accurate, consistent, and aligned with human perception. Looking ahead, such a model might also serve as a reference framework for emerging AI-based document-level evaluations. Current AI systems often provide holistic but qualitative judgments with uncertain completeness; a calibrated non-linear scoring model could, in principle, offer a quantitative reference point for future systems that integrate or adjust their evaluations against established calibrated 0–100 quality scale (user-specific PT/MSV mapping).
@@ 6.1 Operational Guidelines for Quality Managers
One immediate benefit of the non-linear model is its ability to inform and standardize quality expectations across variable evaluation sample sizes. Quality managers can use the calibrated log function to define clear error tolerance thresholds for different content types or use cases, reducing ambiguity when evaluating translation outputs that differ in word count.
For example, a quality manager might specify: “Our reference wordcount is 2000 words with a tolerance of 7 minor errors. For evaluation sample sizes between 1500 and 3000 words, the standard linear approximation is acceptable. Outside that range, the logarithmic model must be used to adjust expectations accordingly.”
This shift allows quality managers and stakeholders to move beyond one-size-fits-all scoring rules and instead adopt quality scores that reflect actual cognitive effort and perceived severity. The result is more fair and interpretable evaluation outcomes, especially when comparing translation outputs across different content types or word counts.
7 Building your non-linear scorecard
We list the following steps in this section including model and least-squares fit, from tolerance to a calibrated score, and scorecard mapping, followed by implications for CAT tools and linguistic quality assessment (LQA) automation, improvement of a unified calibrated quality scale, applications beyond human translation, and document-level evaluation. [^9] The designations and symbols in this section follow the variable naming used in the ASTM working item WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes the MQM scoring models and their mechanics in detail. http://www.astm.org/workitem-wk46396
@@ 7.1 Model and least-squares fit
We model the pass/fail tolerance as a function of sample size $x$ (words) by
@@
Given tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$, estimate $(a,b)$ by least squares
@@
Profiling out $a$ yields, for any fixed $b>0$,
$a(b)\;=\;\frac{\sum_{i}E_{i}\,\ln(1+b\,x_{i})}{\sum_{i}\ln^{2}(1+b\,x_{i})},\qquad\text{then minimize }S(b)=\sum_{i}\!\bigl[E_{i}-a(b)\ln(1+b\,x_{i})\bigr]^{2}\text{ in }b>0,$ (3)
@@
If $x$ was fitted in pages with $W$ words/page and you now compute in words, only $b$ rescales: $b_{\text{words}}=b_{\text{pages}}/W$; $a$ is unchanged.
@@ 7.2 From tolerance to a calibrated score
For an evaluation of size $x=\text{EWC}$ (words), compute the allowed penalty points at the passing threshold
@@
Let $\mathrm{APT}$ be the absolute penalty total from the MQM table. Define the quality fraction
@@
Map $\mathrm{QF}$ to the standard calibrated scale using Maximum Score Value (MSV), Passing Threshold (PT), and Defined Passing Interval $\mathrm{DPI}=\mathrm{MSV}-\mathrm{PT}$:
@@
The rating is Pass if $\mathrm{APT}\leq E_{\text{allowed}}(x)$, else Fail.
Monotonicity and clipping. Since $\mathrm{OS}(x)=\mathrm{PT}+\mathrm{DPI}\bigl(1-\mathrm{APT}/E_{\mathrm{allowed}}(x)\bigr)$ is affine in $\mathrm{APT}$, it decreases monotonically as $\mathrm{APT}$ grows; displaying $\mathrm{OS}_{\rm disp}=\min\{\mathrm{MSV},\max[0,\mathrm{OS}(x)]\}$ does not affect Pass/Fail, which always uses the un‑clipped inequality $\mathrm{APT}\leq E_{\mathrm{allowed}}(x)$.

### v4

@@ 5.3 Anchoring and Calibration Procedures
For operational deployment, we distinguish two adaptation steps:
  • Anchoring: specifying an acceptable error limit at a reference sample size, expressed in severity-weighted penalty points (e.g., “no more than 5 points per 1,000 words”).
  • Calibration: mapping the observed penalty total for a sample into a normalized 0–100 quality scale, supporting interpretability in dashboards and reporting.
Two-point feasibility. Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
@@
Intuitively, the longer sample must permit more errors, but sublinearly, relative to its length increase.
Two-point calibration (numerical). Let $r=E_{1}/E_{0}$ and define $f(b)=\ln(1+bx_{1})-r\,\ln(1+bx_{0})$. Under the feasibility condition above, $f$ has exactly one root $b^{\star}>0$. The parameter $b^{\star}$ can be obtained with a robust one-dimensional root-finding method (e.g., bisection or Brent), after which
@@
See Calibration from Two Points (Numerical) for a one-dimensional root-finding implementation.
@@
For the example anchor points $(x_{0},E_{0})=(1000,5)$ and $(x_{1},E_{1})=(250,2)$, solving $f(b)=0$ yields $b^{\star}\approx 2.880\times 10^{-3}$ and $a^{\star}\approx 3.688$, giving
@@
Fallback for multiple points. If more than two tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$ are available, estimate $(a,b)$ via constrained least squares:
@@
Given any $b_{0}>0$, a fast one-dimensional search over $b$ with closed-form
$a(b)=\frac{\sum_{i}E_{i}\,\ln(1+bx_{i})}{\sum_{i}\ln^{2}(1+bx_{i})}$
provides an efficient initialization (followed by optional joint refinement of $(a,b)$).
For small $b$, $\ln(1+bx)\approx bx$, implying that the model is locally linear near the reference size. This explains why calibrated linear scoring can be adequate only in a neighborhood around its anchor point.
Figure 4: Three interpolations. The spreadsheet trendline $c+k\ln x$ is shown for comparison only; calibration uses $E(x)=a\ln(1+bx)$, which passes through the origin and is anchorable to tolerance points.
5.4 Near-Reference Fidelity Interval for Linear Approximation ($\pm 20\%$)
In practice, a linear rule anchored at a reference size $x_{\mathrm{ref}}$ is often used:
@@ 5.4 Near-Reference Fidelity Interval for Linear Approximation ($\pm 20\%$)
To characterize where this linear approximation remains acceptable, we define the $\pm 20\%$ fidelity interval around $x_{\mathrm{ref}}$ as the set of $x>0$ such that
@@
Writing $\varepsilon=0.20$ and $\alpha=\dfrac{\ln(1+b\,x_{\mathrm{ref}})}{b\,x_{\mathrm{ref}}}$, the boundary points admit a closed form using the Lambert $W$ function:
@@
where $x_{-}$ corresponds to the lower boundary and $x_{+}$ to the upper boundary. [^7] The branch $W_{-1}$ yields the solution near $x_{\mathrm{ref}}$.
Using the illustrative calibration in Section 5.3 (a=3.688,b=0.00288a=3.688,\,b=0.00288), the resulting fidelity intervals are:
  • $\boldsymbol{x_{\mathrm{ref}}=1000}$: $x\in[579,\;1460]$ words.
  • $\boldsymbol{x_{\mathrm{ref}}=2000}$: $x\in[1307,\;2747]$ words.
These intervals quantify the limited range over which a single anchored linear slope remains within 20% relative deviation. Outside this range, proportional extrapolation becomes increasingly inaccurate, motivating use of the logarithmic model or re-anchoring the linear model to a more appropriate reference size.
5.5 Visual comparison across sample-size regimes
@@
Figure 5: Linear model vs. non-linear tolerance. The shaded bands indicate the $\pm 20\%$ fidelity zone around the linear anchor; outside these bands the linear rule under- or over-estimates tolerance.
Figure 5 illustrates that once the slope is fixed by a single anchor point, a linear rule diverges substantially from the concave tolerance pattern implied by human judgments at both smaller and larger evaluation sizes. In contrast, the logarithmic model remains consistent with the length-dependent tolerance trend across the range of sample sizes considered.
@@
Figure 6: Applicability across evaluation sizes. SQC (micro, $\lesssim 250$ words), calibrated linear scoring (meso; reliable only near anchor), and non-linear scoring (macro). Y-axis: tolerance in penalty points.
Finally, Table 1 summarizes operational guidance for selecting scoring models under varying sampling conditions.
@@
EWC varies within $\pm 20\%$ of the anchor Calibrated linear model (re-anchor if the range shifts)
@@
Very short samples ($\lesssim 250$ words) Statistical Quality Control (SQC)
Table 1: Operational guidance for model selection.
@@ 6 Practical Implications
If error tolerance scales non-linearly with sample length—as suggested by the elicitation results in Section 3 and the theoretical considerations in Section 4—then adopting a logarithmic tolerance function has practical implications for translation quality evaluation procedures, quality management policies, and supporting software tooling. In particular, replacing proportional length normalization with a calibrated, length-dependent tolerance model can reduce systematic length-induced bias and improve the consistency of acceptability decisions across samples of varying size.
A length-dependent scoring model may also be relevant for emerging AI-assisted document-level evaluation workflows. Contemporary systems can provide holistic judgments, but their outputs are often qualitative and may vary in completeness or granularity across documents. A calibrated non-linear scoring model offers a quantitative reference function that could be used to normalize or contextualize such judgments against a stable, institution-specific 0–100 scoring scale (e.g., via a user-specific PT/MSV mapping), thereby facilitating comparability across evaluation settings.
@@ 6.1 Operational Guidelines for Quality Managers
A direct operational benefit of the proposed model is support for consistent tolerance setting across variable evaluation sample sizes. Quality managers can use the calibrated logarithmic function $E(x)$ to define explicit acceptability thresholds as a function of evaluation word count, and to adapt tolerances across content types or use cases without relying on ad hoc proportional extrapolation.
For example, a quality policy can be expressed as follows: “We anchor tolerance at 2,000 words with a maximum of 7 minor-error equivalents. Linear approximation is acceptable within the fidelity interval (approximately 1,500–3,000 words in this configuration). Outside this range, tolerance thresholds should be computed using the logarithmic model.”
This approach enables stakeholders to move beyond a single fixed-length scorecard and instead apply acceptance thresholds that reflect length-dependent judgment behavior. The resulting evaluation outcomes are more interpretable and more equitable across texts of different lengths, particularly when quality monitoring requires comparison across heterogeneous content types and evaluation sample sizes.
6.2 Tooling Implications
From a tooling perspective, the proposed model is straightforward to integrate into existing MQM-based evaluation pipelines. Most production systems already compute the severity-weighted penalty total (APT) and the evaluated word count (EWC); therefore, incorporating non-linear length dependence requires only replacing a constant tolerance threshold with a length-dependent tolerance function $E(x)$. Concretely, the acceptability condition becomes
$\mathrm{APT}\leq E(\mathrm{EWC}),$
rather than $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$ under proportional scaling.
Because the core annotation workflow, error typology, and penalty aggregation remain unchanged, the model can be deployed without modifying MQM schemas, annotation tools, or reviewer training protocols. The primary implementation changes occur at the scoring layer and in reporting interfaces (e.g., dashboards and QA summaries), which must display tolerances and pass/fail thresholds that vary with sample length. In addition, calibration parameters $(a,b)$ can be stored per institution, content type, or language pair, enabling consistent application within a unified scoring framework while preserving user-specific acceptance standards.
Implementation sketch.
Algorithm 1 summarizes a minimal integration of length-dependent tolerance into a typical MQM scoring pipeline.
Algorithm 1 Length-dependent tolerance scoring (non-linear MQM)
1:Annotated errors with severities; evaluated word count $\mathrm{EWC}$; calibrated parameters $(a,b)$; score mapping function $\mathrm{Score}(\cdot)$
2:Pass/Fail decision; normalized 0–100 score
3:Compute severity-weighted penalty total $\mathrm{APT}$ from annotations
4:Compute tolerance threshold $T\leftarrow a\ln(1+b\,\mathrm{EWC})$
[…]
7:Compute deviation ratio $d\leftarrow\mathrm{APT}/T$
8:Compute normalized score $s\leftarrow\mathrm{Score}(d)$
9:return $\mathrm{Pass},\,s$
7 Building a Non-Linear Scorecard
This section provides a practical procedure for constructing and deploying a non-linear MQM scorecard, covering (i) model fitting from tolerance points, (ii) conversion from tolerance to a calibrated score on a 0–100 scale, and (iii) scorecard field computation for operational reporting. We further discuss implications for CAT tools and LQA automation, the benefits of a unified calibrated quality scale, applications to AI-generated content, and document-level evaluation considerations. [^8] The designations and symbols in this section follow the variable naming used in the ASTM working item WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes MQM scoring models and scoring mechanics in detail. http://www.astm.org/workitem-wk46396.
@@ 7.1 Model and least-squares fit
We model pass/fail tolerance as a function of evaluation sample size $x$ (in words) using the logarithmic form
@@
Given tolerance points $\{(x_{i},E_{i})\}_{i=1}^{n}$, the parameters $(a,b)$ can be estimated by constrained least squares:
@@
For any fixed $b>0$, the optimal $a$ admits the closed-form expression
$a(b)\;=\;\frac{\sum_{i}E_{i}\,\ln(1+b\,x_{i})}{\sum_{i}\ln^{2}(1+b\,x_{i})}.$ (3)
Substituting $a(b)$ yields a one-dimensional objective
$S(b)=\sum_{i}\!\bigl[E_{i}-a(b)\ln(1+b\,x_{i})\bigr]^{2},$
@@
If the model is fitted using pages with $W$ words per page and later applied using word counts, only $b$ rescales: $b_{\text{words}}=b_{\text{pages}}/W$, while $a$ remains unchanged.
@@ 7.2 From tolerance to a calibrated score
For an evaluation sample of size $x=\mathrm{EWC}$ (words), the tolerance threshold at the passing boundary is computed as
@@
Let $\mathrm{APT}$ denote the observed absolute penalty total from the MQM annotation table. We define the quality fraction as
@@
The quality fraction is then mapped onto a calibrated score using the Maximum Score Value (MSV), Passing Threshold (PT), and Defined Passing Interval $\mathrm{DPI}=\mathrm{MSV}-\mathrm{PT}$:
@@
The evaluation outcome is Pass if $\mathrm{APT}\leq E_{\text{allowed}}(x)$, and Fail otherwise.
Monotonicity and clipping. Since $\mathrm{OS}(x)$ is affine in $\mathrm{APT}$, it decreases monotonically as $\mathrm{APT}$ increases. The displayed score $\mathrm{OS}_{\text{disp}}$ is clipped for reporting convenience and does not affect pass/fail decisions, which are determined solely by the inequality $\mathrm{APT}\leq E_{\text{allowed}}(x)$.